OpenStudy (anonymous):

Which equation represents the line passing through the points (1, 1) and (−2, 13)?

OpenStudy (anonymous):

4x − y = 5 x − 4y = 5 x + 4y = 5 4x + y = 5

OpenStudy (farmdawgnation):

@Bárbara_de_Medeiros Please do not merely provide answers in the future. Our site is about guiding askers to the answers they seek and tutoring them. Not just solving the problem.

OpenStudy (anonymous):

@farmdawgnation so sorry, will do better next time ;D

OpenStudy (anonymous):

So, let me try this again: what you have to do is try to fit those numbers (1,1) and (-2, 13) in the spaces of the x and the y. For example, the first equation: 4x - y = 5 4 . 1 - 1 = 5 4 - 1 = 5 3 = 5. Please note that after doing with the first pairing, you have to do we the second one too. So that's not the answer. If you keep testing each one of them, you'll be able to see that the last one, 4x + y = 5 is the correct, because 4 . 1 + 1 = 5 4 + 1 = 5 5 = 5 And with (-2, 13) 4 . (-2) + 13 = 5 -8 + 13 = 5 5 = 5 So that's the correct answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome! ^^

OpenStudy (farmdawgnation):

Thank you for providing a more detailed answer @Bárbara_de_Medeiros. I won't delete this one, but FYI - in the future it would be ideal to try and convince the asker to guide you through their thought process and/or attempts to work the problem. Cheers.

OpenStudy (anonymous):

Thanks @farmdawgnation.